# Dichography: Two-frame Ultrafast Imaging from a Single Diffraction Pattern

### v1

@@ III.1 Seeing double from a single shot
The first set of results concerns the analysis of diffraction patterns produced by two particles hit by the same XFEL shot, known as “double-hits”. Here, pairs of silver nanoparticles, simultaneously present in the XFEL focus, are intercepted by a single light pulse delivered by SwissFEL [33]. When the particles are far enough apart, the interference term between the two scattered fields is not resolved by the detector. The recorded diffraction image is, in practice, the incoherent sum of the two intensities, suitable for imaging with Dichography. The two retrieved frames correspond to the electron density of the two different particles. Further experimental information are provided in Appendix D.1.
@@ V Outlook
Another modern operational mode of XFELs allows for the generation of light pulses with sub-femtosecond durations [5, 4, 33, 3, 2, 1, 41]. The combination of these attosecond pulses with diffraction imaging has already demonstrated the ability to access ultrafast ionization dynamics by exploiting transient electronic resonances [42]. Recently, attosecond pulses have been combined with the two-color mode [1]. Thanks to the short time duration of the pulses, which can outrun the lifetime of excited electronic configurations, two-color attosecond Dichography may provide unprecedented insights into such elusive states, like the lifetimes, spatial distribution and evolution of electronic excitations or plasmonic phenomena such as charge density waves in complex nanostructures, depending on their size, shape and surrounding environment.
@@ Acknowledgments
The experimental campaigns were conducted at European XFEL (proposal 2857) and SwissFEL (proposal 20222035). We acknowledge the Paul Scherrer Institute, Villigen, Switzerland for provision of free-electron laser beam time at the Maloja instrument of the SwissFEL ATHOS branch and thank the SwissFEL staff for their assistance. We acknowledge the European XFEL in Schenefeld, Germany, for the provision of X-ray free-electron laser beam time at the SQS instrument and thank the EuXFEL staff for their assistance. AC and DR acknowledge the Swiss National Science Foundation (via grant no. 200021E_193642, grant no. 200021-232306, and the NCCR MUST) and ETH Zurich (via collaborative grant 23-2ETH-050) for financial support. TF and CP acknowledge funding by the Deutsche Forschungsgemeinschaft within CRC 1477 ”Light-Matter Interactions at Interfaces” (project number 441234705). We thank the IT Services Group (ISG) of the Department of Physics at ETH Zurich for the excellent support and management of the computing hardware on which the equinox software has been developed and tested.
@@
LH, BB, TB, SDe, ADF, SDo, TF, RH, SK, BK, AØL, BL, SM, TM, MMe, CP, TP, BS, KSi, FSt, RMPT, PT, SU, YO, MMu, DR and AC contributed in the planning and conduction of the experiment at the SQS instrument of the European XFEL (proposal ID 2857). The experiment was led by MMu and YO. LH, AAH, JB, CB, SDo, TF, FG, CG, GJ, MJ, GK, KK, CP, SR, AS, MS, FSc, KSc, ZS, PT, CFU, VW, XX, MY, OY, NY, HZ, BvI, DR and AC contributed in the planning and conduction of the experiment at the Maloja endstation of SwissFEL (proposal ID 20222035). The experiment was led by BvI and DR. SD, JB and SR curated the sample delivery, whose performance allowed for the systematic recording of double-hits. AC developed the original algorithmic approach to Dichography and its numerical implementation. AC and DR analyzed the performance of the method on simulated data and “double-hit” patterns from the SwissFEL experiment. LH pre-processed the two-color data from the European XFEL experiment, took care of the data selection and developed and implemented the size and shape determination of the helium nanodroplets. AC and LH adapted the DCDI method for Dichography and conducted tests on simulated and experimental data. AC, LH and DR contributed in the writing of the manuscript, with input and feedback from all coauthors.
@@ A.2 Constraining experimental data
An intuitive description of the operation is reported in Fig. 6. For a given pixel coordinate $i,j$, the values of the scattered fields $\psi_{ij}^{A}$ and $\psi_{ij}^{B}$ can be represented in the complex plane, as in Fig. 6a, in polar coordinates. The two moduli, $M^{A}$ and $M^{B}$, can then be considered as the two components of a two-dimensional vector, as shown in Fig. 6b. The condition in Eq. (2) can be achieved by a renormalization of the vector $(M^{A},M^{B})$ to obtain a new vector $(M^{\prime A},M^{\prime B})$ whose norm corresponds to the square root of the experimental data, $\sqrt{I}$, while its phase $\phi^{M}$ is preserved. The rescaled moduli $M^{\prime A}$ and $M^{\prime B}$ replace the original ones in the two field estimates, as shown in Fig. 6c, while the field phases $\phi^{A}$ and $\phi^{B}$ are retained. Once the operation is repeated for all entries in the matrix, the experimental constraint is fully back-propagated to the two frames of the reconstruction. A detailed and formal presentation of the mathematical operations involved is provided in Sec. S3 of the Supplemental Material [34].
@@ D.1 Silver nanoparticles at SwissFEL
The two-color experiment was performed at the Maloja endstation of SwissFEL [66, 67]. Wet-chemically grown silver nanoparticles were injected into the experimental chamber, kept under vacuum conditions, using an electrospray and a set of aerodynamic lenses [68, 69]. The electrospray disperses the liquid containing the dissolved nanoparticles into sub-micron-sized droplets. The nanoparticle concentration is tuned to ensure a single sample per droplet. The droplets flow toward the interaction region along with a buffer gas in a set of aerodynamic lenses, where the liquid evaporates.
@@ Appendix E Retrieving the shape of helium droplets
For a diffraction signal from the same pure-helium droplet, composed of the incoherent sum of two different colors, Ref. [35] shows that it is possible to retrieve the size of the droplet and the intensity of the two pulses by fitting the recorded radial intensity profile as the incoherent sum of two Mie-calculated intensities, one for each color. In this procedure, the droplet size $D$ is assumed constant between the two scattering events, while the wavelengths of the two pulses, $\lambda^{A}=1.25,\unit\nano$ and $\lambda^{B}=1.04,\unit\nano$, as well as the corresponding refractive indices for helium, $n^{He}(\lambda^{A})$ and $n^{He}(\lambda^{B})$ [47], are known a priori.

### v2

@@ III.1 Seeing double from a single shot
The first set of results concerns the analysis of diffraction patterns produced by two particles struck by the same XFEL shot, known as “double-hits”. Although they are of lower scientific relevance compared to two-color diffraction images, they serve as a valuable probe for investigating the capabilities, performance, and key features of Dichography.
Here, pairs of silver nanoparticles, simultaneously present in the XFEL focus, are intercepted by a single light pulse delivered by SwissFEL [33]. When the particles are far enough apart, the interference term between the two scattered fields is not resolved by the detector. The recorded diffraction image is, in practice, the incoherent sum of the two intensities, suitable for imaging with Dichography. The two retrieved frames correspond to the electron density of the two different particles. Further experimental information are provided in Appendix D.1.
@@ V Outlook
Another modern operational mode of XFELs allows for the generation of light pulses with sub-femtosecond durations [5, 4, 33, 3, 2, 1, 41]. The combination of these attosecond pulses with diffraction imaging has already demonstrated the ability to access ultrafast ionization dynamics by exploiting transient electronic resonances [42]. Recently, attosecond pulses have been combined with the two-color mode [1]. Thanks to the short time duration of the pulses, which can outrun the lifetime of excited electronic configurations, two-color attosecond Dichography may provide unprecedented insights into such elusive states, like the lifetimes, spatial distribution and evolution of electronic excitations [43, 44] or plasmonic phenomena such as charge density waves in complex nanostructures, depending on their size, shape and surrounding environment.
@@ Acknowledgments
The research leading to these results was supported by SNSF project 10003697 “Visualizing coherent and collective electron dynamics in nanoscale matter” and the COST Action CA21101 “Confined Molecular Systems: From a New Generation of Materials to the Stars (COSY)”. We acknowledge the Paul Scherrer Institute, Villigen, Switzerland for provision of free-electron laser beam time at the Maloja instrument of the SwissFEL ATHOS branch and thank the SwissFEL staff for their assistance. We acknowledge the European XFEL in Schenefeld, Germany, for the provision of X-ray free-electron laser beam time at the SQS instrument and thank the EuXFEL staff for their assistance. TF and CP acknowledge funding by the Deutsche Forschungsgemeinschaft within CRC 1477 ”Light-Matter Interactions at Interfaces” (project number 441234705). We thank the IT Services Group (ISG) of the Department of Physics at ETH Zurich for the excellent support and management of the computing hardware employed for this research.
@@
LH, BB, TB, SDe, ADF, SDo, TF, RH, SK, BK, AØL, BL, SM, TM, MMe, CP, TP, BS, KSi, FSt, RMPT, PT, SU, YO, MMu, DR and AC contributed in the planning and conduction of the experiment at the SQS instrument of the European XFEL (ID 2857). The experiment was led by MMu and YO. LH, AAH, JB, CB, SDo, TF, FG, CG, GJ, MJ, GK, KK, CP, SR, AS, MS, FSc, KSc, ZS, PT, CFU, VW, XX, MY, OY, NY, HZ, BvI, DR and AC contributed in the planning and conduction of the experiment at the Maloja endstation of SwissFEL (ID 20222035). The experiment was led by BvI and DR. SD, JB and SR curated the sample delivery, whose performance allowed for the systematic recording of double-hits. AC developed the original algorithmic approach to Dichography and its numerical implementation. AC and DR analyzed the performance of the method on simulated data and “double-hit” patterns from the SwissFEL experiment. LH pre-processed the two-color data from the European XFEL experiment, took care of the data selection and developed and implemented the size and shape determination of the helium nanodroplets. AC and LH adapted the DCDI method for Dichography and conducted tests on simulated and experimental data. AC, LH and DR contributed in the writing of the manuscript, with input and feedback from all coauthors.
@@ A.2 Constraining experimental data
An intuitive description of the operation is reported in Fig. 6. For a given pixel coordinate $i,j$, the values of the scattered fields $\psi_{ij}^{A}$ and $\psi_{ij}^{B}$ can be represented in the complex plane, as in Fig. 6a, in polar coordinates. The two moduli, $M^{A}$ and $M^{B}$, can then be considered as the two components of a two-dimensional vector, as shown in Fig. 6b. The condition in Eq. (2) can be achieved by a renormalization of the vector $(M^{A},M^{B})$ to obtain a new vector $(M^{\prime A},M^{\prime B})$ whose norm corresponds to the square root of the experimental data, $\sqrt{I}$, while its phase $\phi^{M}$ is preserved. The rescaled moduli $M^{\prime A}$ and $M^{\prime B}$ replace the original ones in the two field estimates, as shown in Fig. 6c, while the field phases $\phi^{A}$ and $\phi^{B}$ are retained. Once the operation is repeated for all entries in the matrix, the experimental constraint is fully back-propagated to the two frames of the reconstruction.
Unlike conventional CDI – in which only a single phase needs to be recovered for each field amplitude – the phase retrieval problem in Dichography is three-fold. Indeed, it requires the recovery of three phase values: $\phi^{A}$ and $\phi^{B}$ for the two scattered fields, as well as $\phi^{M}$, which encodes the relative fields amplitude.
Differently from conventional CDI – where for each field amplitude only as single phase has to be recovered –, the phase retrieval problem in Dichography is three-fold, requiring the retrieval of three phase values, $\phi^{A}$ and $\phi^{B}$ for the two scattered fields and $\phi^{M}$ that encodes their relative amplitude.
A detailed and formal presentation of the mathematical operations involved is provided in Sec. S3 of the Supplemental Material [34].
@@ D.1 Silver nanoparticles at SwissFEL
The two-color experiment was performed at the Maloja endstation of SwissFEL [68, 69]. Wet-chemically grown silver nanoparticles were injected into the experimental chamber, kept under vacuum conditions, using an electrospray and a set of aerodynamic lenses [70, 71]. The electrospray disperses the liquid containing the dissolved nanoparticles into sub-micron-sized droplets. The nanoparticle concentration is optimized for a single sample per droplet. The droplets flow toward the interaction region along with a buffer gas in a set of aerodynamic lenses, where the liquid evaporates.
@@ Appendix E Retrieving the shape of helium droplets
For a diffraction signal from the same pure-helium droplet, composed of the incoherent sum of two different colors, Ref. [35] shows that it is possible to retrieve the size of the droplet and the intensity of the two pulses by fitting the recorded radial intensity profile as the incoherent sum of two Mie-calculated intensities, one for each color. In this procedure, the droplet size $D$ is assumed constant between the two scattering events, while the wavelengths of the two pulses, $\lambda^{A}=1.25\,\unit\nano$ and $\lambda^{B}=1.04\,\unit\nano$, as well as the corresponding refractive indices for helium, $n^{He}(\lambda^{A})$ and $n^{He}(\lambda^{B})$ [49], are known a priori.